OpenStudy (anonymous):

Find b: e^b - e^(1+b) = -4

OpenStudy (anonymous):

think X^a - X^b

OpenStudy (anonymous):

so \[e ^{b-(1+b)}=-4\]

OpenStudy (anonymous):

How did you get that? Could you please show me?

OpenStudy (anonymous):

do you remember the basic exponential rules \[x ^{a}-x ^{b}=x ^{a-b}\]

OpenStudy (anonymous):

just because the base is e does not mean it is any different.

OpenStudy (anonymous):

Now I get it.

OpenStudy (anonymous):

But how can we get the answer for b when there is a negative 4

OpenStudy (anonymous):

Aren't we supposed to log both sides?

OpenStudy (anonymous):

okay here try this as well this would make more sense, try to simplify by taking out e^b from both sides

OpenStudy (anonymous):

so \[e^b-e ^{1+b}=-4 \] becomes \[(1-e)e^b=-4\]

OpenStudy (anonymous):

have x=e^b so to make it easier to see ; think (1-e)x=-4

OpenStudy (anonymous):

do you follow me so far?

OpenStudy (anonymous):

I didn't get this part : (1−e)e^b=−4

OpenStudy (anonymous):

okay I know it looks kind of hard, okay take (1-e)e^b=-4 and expand it what do you get?

OpenStudy (anonymous):

So you factorised

OpenStudy (anonymous):

It's confusing

OpenStudy (anonymous):

Now I get it

OpenStudy (anonymous):

Then what do you do next?

OpenStudy (anonymous):

I think I found the answer

OpenStudy (anonymous):

i know, but trust me its not that hard.

OpenStudy (anonymous):

oh so you get whats your answer?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

0.849?

OpenStudy (anonymous):

that's what i got, so i think you're right! (sorry for late response)

OpenStudy (anonymous):

Thanks for your help. Are you familiar with addition of ordinates?

OpenStudy (anonymous):

Can u give ms an example?